# Anatomical Foundation Models for Brain MRIs

### v1

@@ 3.1.1 Data preparation and preprocessing
All the images underwent the same standard VBM preprocessing, using CAT12 [15], which includes non-linear registration to the MNI template and gray matter (GM) extraction. The final spatial resolution is 1.5mm isotropic and the images are of size 121 x 145 x 121. The preprocessing is performed using the brainprep package  [^1] https://brainprep.readthedocs.io/en/latest/. After preprocessing, we consider in this work modulated gray matter (GM) images, as in [13].
3.2 Background
3.3 Weakly contrastive learning
To tackle the issue of continuous labels, weakly contrastive approaches employ the notion of a degree of “positiveness” between samples. This degree is usually defined by a kernel function $w_{i}=K(y-y_{i})$, where $0\leq w_{i}\leq 1$ is computed by a Gaussian or a Radial Basis Function (RBF) kernel and $y,y_{i}$ are the continuous attributes of interest. The goal of weakly contrastive learning is thus to learn a parametric function $f:\mathcal{X}\rightarrow\mathbb{S}^{d}$ that maps samples with a high degree of positiveness ($w_{i}\sim 1$) close in the latent space and samples with a low degree ($w_{i}\sim 0$) far away from each other. y-Aware [12] proposes to do so by aligning positive samples with a strength proportional to the degree, i.e.:
$\mathcal{L}^{y-aware}=-\sum_{i}\frac{w_{i}}{\sum_{j}w_{j}}\log\left(\frac{\exp(sim(f(x),f(x_{i}))}{\sum_{t=1}^{N}\exp(sim(f(x),f(x_{t})))}\right)$ (1)
where $sim$ is a similarity function (e.g. cosine). A similar approach is also adopted by ExpW [1]. A limitation of this class of approaches is the reliance on one single attribute (age) to determine the alignment strength. They are thus not suited to leverage multiple attributes (e.g. anatomical measurements of the brain). Our proposed approach aims to solve this issue by extending the weakly contrastive formulation to include multiple attributes.
3.4 Proposed method
In this section, we will discuss our approach to include multiple attributes in a weakly contrastive paradigm.
3.4.1 Anatomical Contrastive Learning with anatomical measures
We propose a novel formulation for weakly contrastive learning to employ multiple anatomical attributes derived from the MRI scan, that we call AnatCL. The features we consider are derived from the standard Desikan-Kiliany parcellation [7], based on the information available in the OpenBHB dataset (see Sec. 3.1). Among the anatomical measurements, we consider cortical thickness (CT), gray matter volume (GMV), and surface area (SA) [13]. To include these measurements inside AnatCL, we propose two different formulation, global and local, based on how the similarity across two brain scans is computed.
Local Descriptor
For each of the 68 regions, we obtain a local anatomical descriptor $\psi^{k}\in\mathbb{R}^{3}$ where $k\in\{1,\dots,68\}$ composed by the three anatomical measurements considered for that specific region. In order to compute the degree of positiveness $w_{i}$ across two samples $x$ and $x_{i}$, given their corresponding anatomical descriptors $\Psi=\{\psi^{1},\dots\,\psi^{68}\}$ and $\Psi_{i}=\{\psi^{1}_{i},\dots\,\psi^{68}_{i}\}$ we compute the average cross-region similarity (which we call local descriptor) as:
$w_{i}=\frac{1}{68}\sum_{k=1}^{68}K(\gamma(\psi^{k}),\gamma(\psi^{k}_{i}))$ (2)
where $K$ can either be a similarity function or a kernel as in [12], and $\gamma(x)$ is a normalization that normalizes each component of $x$ to a standard range $[0,1]$. In this work, for simplicity, we choose to employ cosine similarity as $K$.
Global Descriptor
Taking a complementary approach, we now consider global anatomy descriptors $\omega^{k}\in\mathbb{R}^{68}$ for the entire brain with $k\in\{1,2,3\}$, that contain the values across all 68 regions for each anatomical measurement. Similarly to above, in order to compute the degree of positiveness $w_{i}$ between two samples $x$ and $x_{i}$ given their global anatomical descriptors $\Omega=\{\omega^{1},\dots,\omega^{3}\}$ and $\Omega_{i}=\{\omega^{1}_{i},\dots,\omega^{3}_{i}\}$ we compute the cross-measurement similarity (global descriptor) as:
$w_{i}=\frac{1}{3}\sum_{k=1}^{3}K(\omega^{k},\omega^{k}_{i})$ (3)
This time we do not need to standardize $\omega^{k}$ and $\omega^{k}_{i}$ as $K$ is evaluated between features of the same type, so the scale of the value is comparable. Also in this case, for the sake of simplicity, we employ a cosine similarity function for $K$.
Our proposed AnatCL loss function is based on y-Aware [12] and integrates the computed $w_{i}$ (either local or global) as in Eq. 1.
3.4.2 Final objective function
For training, we also consider the available age information, thus our final loss formulation becomes:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{age}+\lambda_{2}\mathcal{L}_{AnatCL}$ (4)
where $\mathcal{L}_{age}$ is defined as in Eq. 1 and $\mathcal{L}_{AnatCL}$ is defined as in Sec. 3.4.1. It is worth noting that the considered anatomical features in AnatCL can be computed directly from the 3D MRI with standard tools such as FreeSurfer  [^2] https://surfer.nmr.mgh.harvard.edu/, and thus our method does not require any additional label in the dataset.
4 Results
In this section, we present the results of our proposed method in comparison to existing state-of-the-art approaches in several downstream task.
@@ 4.1 Experimental setup
Our experiments involved pretraining a model using the proposed AnatCL loss on the OpenBHB dataset, followed by testing the resulting model on various downstream tasks across different datasets. The experimental settings for the two loss formulations, local (AnatCL-G3) and global (AnatCL-L3), are identical. We pretrain two ResNet-18 3D models using VBM-preprocessed images and their corresponding Desikan measures with the proposed formulations. The training process employs the Adam optimizer with a learning rate of 0.0001, and a decay rate of 0.9 applied after every 10 epochs. The models are trained with a batch size of 32 for a total of 300 epochs. As values of $\lambda_{1}$ and $\lambda_{2}$, for simplicity, we use 1. As standard practice in CL approaches [5, 19], the contrastive loss is computed on a fully-connected projection head following the encoder, composed of two layers. To ensure robust evaluation, we perform cross-validation using 5 folds. The results are computed in terms of mean and standard deviation across the 5 folds.
After the pre-training step we evaluate the models by testing their performance with a transfer learning approach: we extract the latent representations generated by the model using only the encoder of the model (i.e., discarding the fully connected head). For each downstream task, we train different linear classifiers on the extracted representations to assess the model’s ability to learn meaningful and generalizable features.
For comparison with standard approaches, we also implement four different baselines: SimCLR [5] (completely self-supervised) and three supervised baselines: a standard model trained with the L1 loss, y-Aware [12] and ExpW [1] which has shown state-of-the-art results on the OpenBHB challenge [13]. For all these methods, we follow the same experimental setup described above.
To run our experiments, we employ a cluster of 4 NVIDIA V100 GPUs, with a single training taking around 10h.
@@ 4.2 Brain age prediction
Preliminary results in terms of brain age prediction and sex classification are evaluated on the OpenBHB dateset. The results are reported in Tab. 2. From the results we conclude that AnatCL is able to match and slightly surpass state-of-the-art performance on brain age prediction. It is worth noting that we do not employ any bias-correction method as in [6, 21]. For sex classification, we do not match ExpW [1], but we improve over the SimCLR and L1 baselines.
Model Method Age MAE Sex
ResNet-18 SimCLR [5] 5.58$\pm$0.53 76.7$\pm$1.67
@@ 4.3 Alzheimer’s desease and cognitive impariments
Model Method HC vs AD sMCI vs pMCI HC vs AD
ResNet-18 SimCLR [5] 78.47$\pm$2.51 61.77$\pm$3.85 73.97$\pm$4.98
@@
In Tab. 3 we report results for Alzheimer’s Disease (AD) detection on ADNI and OASIS-3. While we do not reach the best results overall, AnatCL can usually improve over the self-supervised baseline (SimCLR) and sometimes over either L1, y-Aware or ExpW.
4.4 Schizophrenia and ASD
Schizconnect ABIDE-I
Model Method SCZ. (Broad) SCZ (Strict) Schizoaff. Bipolar Autism Aspergers PDD-NOS
ResNet-18 SimCLR [5] 58.53$\pm$3.52 68.47$\pm$8.47 51.23$\pm$11.94 67.34$\pm$9.96 54.45$\pm$2.99 59.61$\pm$2.72 52.12$\pm$6.62
4.5 Cognitive score / assessments
As final experiments, we turn our attention to predicting clinical assessments score from brain MRIs. To the best of our knowledge, this has not been explored in other works, and it could provide useful insights on the relation between brain anatomy and behavioral phenotypes. The 10 phenotypes considered can be distinguished based on the prediction task: AIMS, depression, handedness and GAIT are classification tasks, while IQ scores (FIQ, VIQ and PIQ) are regression tasks. For handendess, we predict right-handed vs other (left-handed or ambi), for depression we classify between absent vs mild and above, for AIMS we classify between none and minimal vs mild and above, for GAIT between normal vs everything else. For more detailed explanation of the possible value in the considered phenotypes we refer to the official documentation [^3] http://schizconnect.org/documentation#data_dictionaries [^4] http://fcon_1000.projects.nitrc.org/indi/abide/ABIDE_LEGEND_V1.02.pdf.
@@ 4.5 Cognitive score / assessments
Table 5: Assessment scores / phenotypes prediction from brain MRIs.
@@
The results are reported in Tab. 5. While we cannot conclude that any of the analysed method can accurately predict all clinical assessments from MRI scans, AnatCL overall achieves the best results three out of ten times, which is more than any other baseline. Interestingly, AnatCL can better predict the overall AIMS score and patients’ handedness, hinting that brain anatomy may be linked with these phenotypes.
5 Limitations, impact and ethical considerations
We believe that foundation models for neuroimaging may have a considerable impact on accurately diagnosing neurological and psyhicatric diseases. With AnatCL we aim at laying the foundations for this path. Currently, AnatCL is limited towards using a single data modality (structural MRI), considering limited anatomical features (only CT, GMV, and SA) and to a relatively small backbone (ResNet-18). Future research should focus on improving these issues, in order to obtain even more accurate predictions. Altough, deep learning techniques can be used in a variety of contexts, we do not believe that AnatCL inherently poses any ethical issue. Furthermore, all the data employed in this work is publicly available for researchers.
6 Conclusions
Appendix A Anatomical Features
A.1 Details on preprocessing
Anatomical measurements on OpenBHB are released as part of the original dataset. All details can be found in [13]. For preprocessing, the brainprep [^5] https://github.com/neurospin-deepinsight/brainprep module was employed, with FreeSurfer version FSL-6.0.5.1, and Desikan-Killiany atlas as provided in FreeSurfer [^6] https://surfer.nmr.mgh.harvard.edu/fswiki/CorticalParcellation.
A.2 Choice of Anatomical Features
OpenBHB provides 7 different anatomical measures, computed with FreeSurfer: Cortical Thickness (CT, mean and std values), Gray Matter Volume (GMV), Surface Area, Integrated Mean, Gaussian Curvature Index, and Intrinsic Curvature Index. In order to select the most relevant features, we conduct a preliminary study by training a simple linear model (Ridge regression) over each feature, with the task of brain age prediction on the OpenBHB data. The results are evaluated in terms of negative mean absolute error and $R^{2}$ score, and are reported in Tab. 6. The best results are achieved by GVM and Surface Area, which are morphological features regarding the volume and the total area of each ROI. Thus, we select these two features for our experiments with AnatCL. In addition to GMV and Surface Area measures, we also include CT as it is a widely used measure in related works [27, 28]. The final choice is thus composed of three features, which cover different aspects of the morphology and geometry of brain ROIs (e.g. cortical thickness is relevant for conditions such as Alzheimer’s Disease [24], and volume and surface measurements can signal brain atrophy, linked to different conditions).
CT (mean) GMV Surf. Area Int. Mean Gauss. Curv. Index Intr. Curv. Index

### v3

@@ 4.1.1 Data preparation and preprocessing
All the images underwent the same standard VBM preprocessing, using CAT12 [16], which includes non-linear registration to the MNI template and gray matter (GM) extraction. The final spatial resolution is 1.5mm isotropic and the images are of size 121 x 145 x 121. The preprocessing is performed using the brainprep package  [^2] https://brainprep.readthedocs.io/en/latest/. After preprocessing, we consider modulated GM images, as in [13].
4.1.2 Computation of anatomical features
Anatomical measurements on OpenBHB are released as part of the original dataset. All details can be found in [13]. For deriving the measuremets, the brainprep [^3] https://github.com/neurospin-deepinsight/brainprep module was employed, with FreeSurfer version FSL-6.0.5.1, and Desikan-Killiany atlas as provided in FreeSurfer [^4] https://surfer.nmr.mgh.harvard.edu/fswiki/CorticalParcellation.
CT (mean) GMV Surf. Area Int. Mean Gauss. Curv. Index Intr. Curv. Index
@@ 4.2 Experimental setup
Our experiments involved pre-training a model using the proposed AnatCL loss on the OpenBHB dataset, followed by testing on various downstream tasks across different datasets. The experimental settings for the two loss formulations, local (AnatCL-G3) and global (AnatCL-L3), are identical. We pre-train two ResNet-18 3D models using VBM-preprocessed images and their corresponding Desikan measures with the proposed formulations. The training process employs the Adam optimizer with a learning rate of 0.0001, and a decay rate of 0.9 applied after every 10 epochs. The models are trained with a batch size of 32 for a total of 300 epochs. As values of $\lambda_{1}$ and $\lambda_{2}$, for simplicity, we use 1. As standard practice in CL approaches [5, 21], the contrastive loss is computed on a fully-connected projection head following the encoder, composed of two layers. To ensure robust evaluation, we perform cross-validation using 5 folds. The results are computed in terms of mean and standard deviation across the 5 folds. After the pre-training step we evaluate the models by testing their performance with a transfer learning approach: we extract the latent representations generated by the model using only the encoder of the model (i.e., discarding the fully connected head). For each downstream task, we train different linear classifiers on the extracted representations to assess the model’s ability to learn meaningful and generalizable features. To run our experiments, we employ a cluster of 4 NVIDIA V100 GPUs, with a single training taking around 10h.
4.2.1 Choice of Anatomical Features
OpenBHB provides 7 different anatomical measures, computed with FreeSurfer: Cortical Thickness (CT, mean and std values), Gray Matter Volume (GMV), Surface Area, Integrated Mean, Gaussian Curvature Index, and Intrinsic Curvature Index. To motivate the choice of the anatomical features, we conduct a preliminary study by training a simple linear model (Ridge regression) over each feature, with the task of brain age prediction on the OpenBHB data. The results are evaluated in terms of negative mean absolute error and $R^{2}$ score and are reported in Tab. 2. The best results are achieved by GVM and Surface Area, which are morphological features regarding the volume and the total area of each ROI. Thus, we select these two features for our experiments with AnatCL. In addition to GMV and Surface Area measures, we also include CT as it is a widely used measure in related works [29, 30]. The final choice is thus composed of three features, which cover different aspects of the morphology and geometry of brain ROIs (e.g. cortical thickness is relevant for conditions such as Alzheimer’s Disease [26], and volume and surface measurements can signal brain atrophy, linked to different conditions).
4.2.2 Baselines for comparison
In this work we focus on contrastive learning has it has been demonstrated to work better than non-contrastive method in recent literature [14]. For this reason, we pick recent state-of-the-art contrastive methods in neuroimaging, namely y-Aware [12] and ExpW [1]. In addition to these recent techniques, we also consider the popular SimCLR framework [5] as self-supervised contrastive baseline. For comparison with traditional non-contrastive supervised techniques, we also compare to baseline models trained with L1 loss for brain age prediction [20].
@@ 4.3 Brain age prediction
The first two downstream tasks we evaluate are brain age prediction and sex classification, on the OpenBHB dataset. The results are reported in Tab. 3. From the results, we conclude that AnatCL is able to slightly surpass state-of-the-art performance on brain age prediction, reaching the lowest absolute error of 2.61 years. It is worth noting that we do not employ any bias-correction method as in [6, 23], but we consider the raw predictions of the model. For sex classification, we do not match ExpW [1], but we improve over the SimCLR and L1 baselines.
Method Age MAE Sex
SimCLR [5] 5.58$\pm$0.53 76.7$\pm$1.67
@@
Schizconnect ABIDE-I
Method SCZ. (Broad) SCZ (Strict) Schizoaff. Bipolar Autism Aspergers PDD-NOS
SimCLR [5] 58.53$\pm$3.52 68.47$\pm$8.47 51.23$\pm$11.94 67.34$\pm$9.96 54.45$\pm$2.99 59.61$\pm$2.72 52.12$\pm$6.62
4.4 Schizophrenia and ASD
@@ 4.5 Alzheimer’s desease and cognitive impariments
Method HC vs AD sMCI vs pMCI HC vs AD
SimCLR [5] 78.47$\pm$2.51 61.77$\pm$3.85 73.97$\pm$4.98
@@
In Tab. 5 we report results for Alzheimer’s Disease (AD) detection on ADNI and OASIS-3. While we do not reach the best results overall, AnatCL can consistenly improve over the self-supervised baseline SimCLR and over the weakly supervised method y-Aware and reaches overall competitive results. On OASIS-3, AnatCL shows also better results than ExpW, in both local and global variants. Overall, AnatCL provides more consistent performance than other methods.
4.6 Clinical assessments and phenotypes
(a) ABIDE-I
(b) SchizConnect
@@ 4.6 Clinical assessments and phenotypes
Figure 4: Assessment scores and phenotypes prediction from brain MRIs.
As final experiments, we turn our attention to predicting clinical assessments score from brain MRIs. To the best of our knowledge, this has not been explored in other works, and it could provide useful insights on the relation between brain anatomy and behavioral phenotypes. The 10 phenotypes considered can be distinguished based on the prediction task: AIMS, depression, handedness and GAIT are classification tasks, while IQ scores (FIQ, VIQ and PIQ) are regression tasks. For handendess, we predict right-handed vs other (left-handed or ambi), for depression we classify between absent vs mild and above, for AIMS we classify between none and minimal vs mild and above, for GAIT between normal vs everything else. For more detailed explanation of the possible value in the considered phenotypes we refer to the official documentation [^5] http://schizconnect.org/documentation#data_dictionaries [^6] http://fcon_1000.projects.nitrc.org/indi/abide/ABIDE_LEGEND_V1.02.pdf. The results are reported in Fig. 4. While we cannot conclude that any of the analysed method can accurately predict all clinical assessments from MRI scans, AnatCL overall achieves the best results three out of ten times, which is more than any other baseline. Interestingly, AnatCL can better predict the overall AIMS score and patients’ handedness, hinting that brain anatomy may be linked with these phenotypes. The link between brain anatomy and involuntary movements could lead to the discovery of novel biomarkers and should be studied more in depth in future research.
4.7 Ablation study
In this section, we perform an ablation study to motivate the design choices we made for AnatCL. Our ablation study aims to assess the benefit of different network architectures (i.e. larger models), and different formulations of the AnatCL loss function.
Model size and architecture
We perform an ablation on different neural architectures, considering larger models such as 3D ResNet-50. We report the results of this ablation on the Schizconnect dataset here in Tab. 6. From the obtained results, we do not find significant advantages in using a 3D ResNet-50 over the simpler and smaller version 3D ResNet-18. For this reason, we chose to employ this simpler architecture, as it is practically more convenient to train and deploy in real settings. It is worth mentioning that similar findings are reported in related literature [13], showing that deeper architectures such as even a 3D DenseNet-121 do not bring significant improvements over a 3D ResNet-18, at least on VBM-processed images. We did not explore different and larger architectures such as ViT [11], as we were limited by computational constraints. In fact, a 3D ViT implementation requires significantly more memory and computational power, which makes it less practical and favorable with respect to ResNet-18, also from a carbon footprint point of view.
Model Method SCZ. (Broad) SCZ. (Strict) Schizoaff. Bipolar
ResNet-18 y-Aware [12] 66.20$\pm$4.50 71.04$\pm$2.31 70.29$\pm$14.73 63.95$\pm$19.81
ExpW [1] 69.53$\pm$4.43 67.65$\pm$8.27 63.26$\pm$18.06 74.34$\pm$18.95
AnatCL-G3 72.55$\pm$5.16 71.03$\pm$8.53 73.65$\pm$7.29 63.43$\pm$15.94
AnatCL-L3 66.38$\pm$5.96 72.07$\pm$8.42 68.37$\pm$8.60 56.60$\pm$11.48
ResNet-50 y-Aware 66.14$\pm$9.23 68.17$\pm$5.45 64.94$\pm$17.13 69.16$\pm$22.00
ExpW 66.39$\pm$6.54 73.69$\pm$5.23 66.07$\pm$14.89 66.19$\pm$19.95
AnatCL-G3 68.07$\pm$4.16 72.16$\pm$5.53 57.92$\pm$14.50 78.71$\pm$18.67
AnatCL-L3 65.20$\pm$3.19 62.99$\pm$8.39 69.40$\pm$9.98 67.61$\pm$15.81
Table 6: Model size ablation study. Larger architectures do not result in consistent improvements over simpler models, as also shown in [13].
Loss formulation and pretraining
We also experiment with different variants of the AnatCL loss. First, we set $\lambda_{2}=0$, meaning that we only consider anatomical features in the training process. We call this variant AnatSSL (as in Self-Supervised), as it effectively does not require any other information besides the anatomical data itself. This variant may be useful when pretraining on a large cohort in which the patient’s age is unknown. We also perform a simple pre-training of the network using an L1 loss to predict the anatomical features:
$\mathcal{L}_{anat-sup}=||r(f(x_{i}))-\bar{\omega}_{i}||$ (7)
where $r:\mathbb{R}^{d}\rightarrow\mathbb{R}^{3}$ is a linear regression layer, and $\bar{\omega}_{i}$ is the average of the patient’s global descriptors:
$\bar{\omega}_{i}=\left[\frac{1}{68}\sum_{j}\omega^{1}_{i,j},\dots,\frac{1}{68}\sum_{j}\omega^{3}_{i,j}\right]$ (8)
We consider the average global descriptor composed of 3 values for this formulation, instead of the average local vector which would contain 68 elements for regression for practical reasons. The results are presented in Tab. 7.
ADNI OASIS-3
Method HC vs AD sMCI vs pMCI HC vs AD
L1 (anat. supervised) 64.63 60.79 67.01
AnatSSL-G3 73.21 60.69 69.77
AnatSSL-L3 76.32 61.11 69.12
AnatCL-G3 80.47 66.03 75.59
AnatCL-L3 80.11 62.83 75.88
Table 7: Loss function and pre-training strategy ablation. AnatCL learns more generalizable representations by combining anatomical features and patients’ metadata (i.e. age).
As we observe from the results, employing only the anatomical features is not enough to obtain more robust representations. However we can still notice how contrastive-based methods such as AnatSSL outperforms traditional supervised baselines. Our proposed approach, on the other hand, leverages information from both the patient and brain anatomy, leading to more robust and generalizable representations.
5 Conclusions
Supplementary Material
6 Limitations, impact, and ethics
We believe that foundation models for neuroimaging may have a considerable impact on accurately diagnosing neurological and psychiatric diseases. With AnatCL we aim at laying the foundations for this path. Currently, AnatCL is limited to using a single data modality (structural MRI), considering limited anatomical features (only CT, GMV, and SA) and to a relatively small backbone (ResNet-18). Future research should focus on improving these issues, in order to obtain even more accurate predictions. Although deep learning techniques can be used in a variety of contexts, we do not believe that AnatCL inherently poses any ethical issue. Furthermore, all the data employed in this work is publicly available to researchers.
7 Results
7.1 Assessment scores and phenotypes
Full numerical values are presented in Tab. 8.
@@ 7.1 Assessment scores and phenotypes
Table 8: Assessment scores / phenotypes prediction from brain MRIs.